sphenix-tpc-l AT lists.bnl.gov
Subject: Sphenix-tpc-l mailing list
List archive
Re: [[Sphenix-tpc-l] ] TPC Meeting TOMORROW MONDAY 03.03 at 9:30 AM ET
- From: "Hughes, Charles [PHYSA]" <chughes2 AT IASTATE.EDU>
- To: Takao Sakaguchi <takao AT bnl.gov>, "sphenix-tpc-l AT lists.bnl.gov" <sphenix-tpc-l AT lists.bnl.gov>, "divaldo95 AT gmail.com" <divaldo95 AT gmail.com>, Gabor David <gabor.david AT stonybrook.edu>, "balazs.ujvari AT cern.ch" <balazs.ujvari AT cern.ch>, "Finnelli, Kristina" <kfinnelli AT bnl.gov>, "Kuczewski, John" <jkuczewski AT bnl.gov>, "Miraglia, Damon" <dmiraglia AT bnl.gov>, "Mandracchia, Kevin" <kmandracchia AT bnl.gov>, shana Prifte <sprifte AT bnl.gov>, Stephen Boose <boose AT bnl.gov>, "Robert P. Pisani" <pisani AT bnl.gov>, Seth Howell <seth.howell AT stonybrook.edu>, "Azmoun, Bob" <azmoun AT bnl.gov>, "Vasquez, Joel" <jvasquez AT bnl.gov>, Frank Toldo <fatoldo AT bnl.gov>, Michael Lenz <mlenz AT bnl.gov>, Thomas K Hemmick <tkhemmick AT gmail.com>, "Huang, Jin" <jhuang AT bnl.gov>, Thomas K Hemmick <Thomas.Hemmick AT stonybrook.edu>, "Haggerty, John" <haggerty AT bnl.gov>, "josborn1 AT bnl.gov" <josborn1 AT bnl.gov>, Luke Legnosky <luke.legnosky AT stonybrook.edu>
- Subject: Re: [[Sphenix-tpc-l] ] TPC Meeting TOMORROW MONDAY 03.03 at 9:30 AM ET
- Date: Mon, 3 Mar 2025 00:18:37 +0000
Hi
@Takao Sakaguchi
Thans a lot for the detailed explanations as well as options for what could be done. For reference, this is the current situation we have in TPC today (you can see the wandering pedestals in many places - at least 1 FEE has it in many TPC sectors):

Lets discuss tomorrow morning.
Thans a lot for the detailed explanations as well as options for what could be done. For reference, this is the current situation we have in TPC today (you can see the wandering pedestals in many places - at least 1 FEE has it in many TPC sectors):

Lets discuss tomorrow morning.
From: Takao Sakaguchi <takao AT bnl.gov>
Sent: Sunday, March 2, 2025 4:56 PM
To: Hughes, Charles [PHYSA] <chughes2 AT IASTATE.EDU>; sphenix-tpc-l AT lists.bnl.gov <sphenix-tpc-l AT lists.bnl.gov>; divaldo95 AT gmail.com <divaldo95 AT gmail.com>; Gabor David <gabor.david AT stonybrook.edu>; balazs.ujvari AT cern.ch <balazs.ujvari AT cern.ch>; Finnelli, Kristina <kfinnelli AT bnl.gov>; Kuczewski, John <jkuczewski AT bnl.gov>; Miraglia, Damon <dmiraglia AT bnl.gov>; Mandracchia, Kevin <kmandracchia AT bnl.gov>; shana Prifte <sprifte AT bnl.gov>; Stephen Boose <boose AT bnl.gov>; Robert P. Pisani <pisani AT bnl.gov>; Seth Howell <seth.howell AT stonybrook.edu>; Azmoun, Bob <azmoun AT bnl.gov>; Vasquez, Joel <jvasquez AT bnl.gov>; Frank Toldo <fatoldo AT bnl.gov>; Michael Lenz <mlenz AT bnl.gov>; Thomas K Hemmick <tkhemmick AT gmail.com>; Huang, Jin <jhuang AT bnl.gov>; Thomas K Hemmick <Thomas.Hemmick AT stonybrook.edu>; Haggerty, John <haggerty AT bnl.gov>; josborn1 AT bnl.gov <josborn1 AT bnl.gov>; Luke Legnosky <luke.legnosky AT stonybrook.edu>
Subject: Re: TPC Meeting TOMORROW MONDAY 03.03 at 9:30 AM ET
Sent: Sunday, March 2, 2025 4:56 PM
To: Hughes, Charles [PHYSA] <chughes2 AT IASTATE.EDU>; sphenix-tpc-l AT lists.bnl.gov <sphenix-tpc-l AT lists.bnl.gov>; divaldo95 AT gmail.com <divaldo95 AT gmail.com>; Gabor David <gabor.david AT stonybrook.edu>; balazs.ujvari AT cern.ch <balazs.ujvari AT cern.ch>; Finnelli, Kristina <kfinnelli AT bnl.gov>; Kuczewski, John <jkuczewski AT bnl.gov>; Miraglia, Damon <dmiraglia AT bnl.gov>; Mandracchia, Kevin <kmandracchia AT bnl.gov>; shana Prifte <sprifte AT bnl.gov>; Stephen Boose <boose AT bnl.gov>; Robert P. Pisani <pisani AT bnl.gov>; Seth Howell <seth.howell AT stonybrook.edu>; Azmoun, Bob <azmoun AT bnl.gov>; Vasquez, Joel <jvasquez AT bnl.gov>; Frank Toldo <fatoldo AT bnl.gov>; Michael Lenz <mlenz AT bnl.gov>; Thomas K Hemmick <tkhemmick AT gmail.com>; Huang, Jin <jhuang AT bnl.gov>; Thomas K Hemmick <Thomas.Hemmick AT stonybrook.edu>; Haggerty, John <haggerty AT bnl.gov>; josborn1 AT bnl.gov <josborn1 AT bnl.gov>; Luke Legnosky <luke.legnosky AT stonybrook.edu>
Subject: Re: TPC Meeting TOMORROW MONDAY 03.03 at 9:30 AM ET
Hi Charles and all,
Thanks for the list of activity summary.
About the wandering pedestals, this is basically due to failed configuration
of the SAMPA chip. We tried to fix the resistors on the I2C for as many FEEs
as possible (around 220 out of 624) in Dec 2023, but about the 400 were
left as they are because the symptom was not seen for them at that time,
and in addition, we didn't have enough time to fix all of them.
If the configuration was properly done, the pedestal values are fixed to 60.
This is realized by the following procedure in the SAMPA:
1, Turning on so-called Baseline-correction-3 (BC3). The BC3 function tries
to follow the long-term perturbation of the baseline values and force them
to zero. This results in the pedestal values of zero, meaning the any signal
going below zero will be cut out and not seen.
2, In order to keep the data below mean pedestal values, we added an offset
of 60 after the BC3 was applied. This brings the default pedestal values to 60.
If the SAMPA configuration was failed, both or either of 1 and 2 are not applied,
and the result will be:
- if 1 is failed, the pedestal value will be ~150. The out-of-box typical pedestal
value is ~80, and if 1 is not applied, it won't be set to zero first, but only offset
is added, meaning the pedestal will be ~150.
- if 2 is failed, the pedestal values stick to zero and you won't see other half of
the Gaussian distribution.
The safest and sure solution to avoid this problem is to apply neither 1 nor 2.
Then you will know the pedestal values are for sure around 80. The downside
of this is that you have to set the zero-suppression threshold channel-by-channel
if you want to set the thresholds in terms of n * RMS rigorously.
But alternately, we could average the threshold values over one SAMPA which we
are doing now anyway, assuming the pedestal mean and RMS values don't change
so much over a SAMPA, which is usually the case.
Shall we do this? It is very easy to do.
By the way, I don't think I will replace any other FEEs now. The only thing I would
like to do is to remove the GND braids from a FEE in the 6 o'clock R3 in south as
we learned this won't solve the noise issue, and the way the GND braids were put
is ugly.
Takao
Thanks for the list of activity summary.
About the wandering pedestals, this is basically due to failed configuration
of the SAMPA chip. We tried to fix the resistors on the I2C for as many FEEs
as possible (around 220 out of 624) in Dec 2023, but about the 400 were
left as they are because the symptom was not seen for them at that time,
and in addition, we didn't have enough time to fix all of them.
If the configuration was properly done, the pedestal values are fixed to 60.
This is realized by the following procedure in the SAMPA:
1, Turning on so-called Baseline-correction-3 (BC3). The BC3 function tries
to follow the long-term perturbation of the baseline values and force them
to zero. This results in the pedestal values of zero, meaning the any signal
going below zero will be cut out and not seen.
2, In order to keep the data below mean pedestal values, we added an offset
of 60 after the BC3 was applied. This brings the default pedestal values to 60.
If the SAMPA configuration was failed, both or either of 1 and 2 are not applied,
and the result will be:
- if 1 is failed, the pedestal value will be ~150. The out-of-box typical pedestal
value is ~80, and if 1 is not applied, it won't be set to zero first, but only offset
is added, meaning the pedestal will be ~150.
- if 2 is failed, the pedestal values stick to zero and you won't see other half of
the Gaussian distribution.
The safest and sure solution to avoid this problem is to apply neither 1 nor 2.
Then you will know the pedestal values are for sure around 80. The downside
of this is that you have to set the zero-suppression threshold channel-by-channel
if you want to set the thresholds in terms of n * RMS rigorously.
But alternately, we could average the threshold values over one SAMPA which we
are doing now anyway, assuming the pedestal mean and RMS values don't change
so much over a SAMPA, which is usually the case.
Shall we do this? It is very easy to do.
By the way, I don't think I will replace any other FEEs now. The only thing I would
like to do is to remove the GND braids from a FEE in the 6 o'clock R3 in south as
we learned this won't solve the noise issue, and the way the GND braids were put
is ugly.
Takao
On 3/2/25 3:00 PM, Hughes, Charles [PHYSA] wrote:
Hello All
Today I'm using the sphenix-tpc list at the suggestion of @Huang, Jin. I'm not sure who all is on that list, so I included some of the Debrecen and BNL folks as well (apologies if you got 2 emails).
Let's meet MONDAY 03.03 at 9:30 AM ET (3:30 PM CET) to discuss ongoing operations/analysis/planning for the TPC. In particular:
HV: Last week we made a lot of progress which has culminated in the TPC running every night and continuously throughout the weekend. However there is still more progress to be made including HV channel mapping and further work on the GUI (implementing weather compensation). Additionally, @Thomas K Hemmick asked me to investigate the cable for NS 12 o'clock R3 and we should discuss plan about that.
Gas/Cooling: @Robert P. Pisanihas helped get our cooling to the point where it has been proven stable for many days. Rob has also requested s few hours when the HV/LV can be off for him to commission some new temperature sensors for the CDU in the hall which will give a much better picture of what's going on there. We should also ask him when the dry cabinet will be installed.
LV/FEEs: Last week we also made progress reducing the noise in the TPC, discovering the EMCAL rack communications module as a big contributor. @Haggerty, John and @Stephen Boose will need to get together to discuss workable solutions. Also @Sakaguchi, Takaohad expressed a desire to explore replacing some FEEs which are newly malfunctioning. Tom had some questions for the Takao and Jin about the TPC FEE settings (why is pedestal set to 60, how to fix the wandering pedestal issue). I would like to talk to Jin and Takao about re-calculation of the TPC ZS thresholds. Since the noise was reduced, last years thresholds are not optimal.
Diffuse Laser: Friday, @Huang, Jin asked if he could re-start services on the diffuse laser fan out board. We asked him to hold off until tomorrow. I think we should bring back the diffuse laser entirely, including some standalone runs to send to Ben Kimmelman for re-analysis. Let's discuss.
Line Laser: @Seth Howell, @Azmoun, Bob, @Finnelli, Kristina, and @Kuczewski, John made great progress on the TPC line laser. This culminated in the testing of the lasers in the NS bore on Friday. We should discuss this as well as @divaldo95 AT gmail.com's work on the noise mitigation. We should also discuss the path forward including Seth's alignment parameter search and getting to the point where we can think about threading the laser into the TPC. Contingency Box: As @Mandracchia, Kevin said, we need to discuss with John when we can ramp up work with the contingency boxes. We should also make sure we take advantage of @shana Prifte's offer to help. Analysis: Now that we have a good amount of cosmic data, we should discuss analysis of this for gain balancing. @josborn1 AT bnl.gov and @Luke Legnosky have started to look around in the data. What do we need to do to get to MIP peaks ? Any other pressing business (directives from the workfest, offline analysis, etc...). For example, firmware upgrade.
Indico Link:https://indico.bnl.gov/event/27062/
Zoom is the leader in modern enterprise video communications, with an easy, reliable cloud platform for video and audio conferencing, chat, and webinars across mobile, desktop, and room systems. Zoom Rooms is the original software-based conference room solution used around the world in board, conference, huddle, and training rooms, as well as executive offices and classrooms. Founded in 2011, Zoom helps businesses and organizations bring their teams together in a frictionless environment to get more done. Zoom is a publicly traded company headquartered in San Jose, CA.urldefense.com
-
[[Sphenix-tpc-l] ] TPC Meeting TOMORROW MONDAY 03.03 at 9:30 AM ET,
Hughes, Charles [PHYSA], 03/02/2025
-
Re: [[Sphenix-tpc-l] ] TPC Meeting TOMORROW MONDAY 03.03 at 9:30 AM ET,
Takao Sakaguchi, 03/02/2025
- Re: [[Sphenix-tpc-l] ] TPC Meeting TOMORROW MONDAY 03.03 at 9:30 AM ET, Hughes, Charles [PHYSA], 03/02/2025
-
Re: [[Sphenix-tpc-l] ] TPC Meeting TOMORROW MONDAY 03.03 at 9:30 AM ET,
Takao Sakaguchi, 03/02/2025
Archive powered by MHonArc 2.6.24.